How many years will the following take? $644 to grow to $13,342 if invested at 16.81 percent, compounded annually.
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

13,342=644*(1.1681)^n

(13,342/644)=(1.1681)^n

Taking log on both sides;

log (13,342/644)=n*log (1.1681)

n=log (13,342/644)/log (1.1681)

=19.51 years(Approx).